# On the multiplier spectrum of polynomials

### v1

@@ 4.3. Ritt move
(4.12) $Q^{\circ n}(z^{k})-z^{k}=(P^{\circ n}(z))^{k}-z^{k}=(P^{\circ n}(z)-z)\prod_{j=1}^{k-1}(P^{\circ n}(z)-\zeta_{k}^{j}z).$
The left-hand side of (4.12) can be regarded as a polynomial in the variable $z^{k}-z_{0}^{k}$, and the right-hand side as a polynomial in $z-z_{0}$. Since $z_{0}\neq 0$, we have ${\rm ord}_{z-z_{0}}(z^{k}-z_{0}^{k})=1$. Applying $\rm ord$} to both sides of (4.12), we obtain
${\rm ord}_{z-z_{0}^{k}}(Q^{\circ n}(z)-z)$ $={\rm ord}_{z^{k}-z_{0}^{k}}(Q^{\circ n}(z^{k})-z^{k})\cdot 1$
$={\rm ord}_{z^{k}-z_{0}^{k}}(Q^{\circ n}(z^{k})-z^{k})\cdot{\rm ord}_{z-z_{0}}(z^{k}-z_{0}^{k})$
$={\rm ord}_{z-z_{0}}(Q^{\circ n}(z^{k})-z^{k})$
$={\rm ord}_{z-z_{0}}(P^{\circ n}(z)-z).$
Thus ${\rm multi}_{F_{1}}(z_{0})={\rm multi}_{F_{2}}(z_{0}^{k})\geq 2$.
For $z_{0}=0=z_{0}^{k}$, since $r\geq 2$, 0 is a super-attracting fixed point of both $P^{\circ n}$ and $Q^{\circ n}$, i.e., $\rho_{P^{\circ n}}(0)=\rho_{Q^{\circ n}}(0)=0$. Hence ${\rm multi}_{F_{1}}(0)={\rm multi}_{F_{2}}(0)=1$, concluding the proof of (4.10).
From (4.10), if $h(z_{0})=h(z_{1})$ implies $z_{0}=z_{1}$ for $z_{0},z_{1}\in F_{1}$ (in which case we say $n$ is “good”), then $S_{n}(P)=S_{n}(Q)$. We show that $n$ is good for all $n$ in an infinite arithmetic progression.
Assume $n\in\mathbb{Z}_{>0}$ is not good. Then there exist $z_{0}\neq z_{1}$ in $F_{1}$ with $z_{1}^{k}=z_{0}^{k}$. Hence $z_{1}\neq 0$ and $z_{1}=\zeta z_{0}$ for some $\zeta\in\mathbb{C}$ with $\zeta^{k}=1$ and $\zeta\neq 1$. It is easy to see that there exists $R_{n}(z)\in\mathbb{C}[z]$ with $R_{n}(0)\neq 0$ such that $P^{\circ n}(z)=z^{r^{n}}\cdot R_{n}(z^{k})$ and $Q^{\circ n}(z)=z^{r^{n}}\cdot R_{n}(z)^{k}$. From $P^{\circ n}(z_{0})=z_{0}$ and $P^{\circ n}(\zeta z_{0})=\zeta z_{0}$, we deduce $\zeta^{r^{n}}=\zeta$, so
$1<{\rm ord}_{\mathbb{C}^{\times}}(\zeta)\mid\gcd(r^{n}-1,k).$
Thus, for every integer $n\geq 1$, either $n$ is good, or $\gcd(r^{n}-1,k)>1$.
Let $M$ be the minimal positive integer such that $\gcd(r(r^{M}-1),k)=1$. In particular, $r$ and $k$ are coprime. By Euler’s theorem, $r^{\varphi(k)}\equiv 1\pmod{k}$. Let $d$ be the minimal positive integer with $r^{d}\equiv 1\pmod{k}$. Then $d\mid\varphi(k)$. For every integer $N\geq 0$, we have
$r^{M+Nd}-1\equiv r^{M}-1\pmod{k}$,
so $\gcd(r^{M+Nd}-1,k)=1$. Therefore, every $n$ in the arithmetic progression $(M+Nd)_{N=0}^{\infty}$ is good. Since the sequence $(r^{n}\bmod k)_{n}$ has exact period $d$, $M\leq d$ by minimality. In fact, we must have $M<d$ (otherwise $k\mid\gcd(r^{d}-1,k)=\gcd(r^{M}-1,k)=1$, contradicting $k>1$). ∎
@@
Given integers $n\geq 1$ and $k\geq 2$, let $V(z)\in\mathbb{C}[z]$ be a polynomial of degree $n$ with $V(0)\neq 0$. Consider the Ritt move
$(P(z):=z\cdot V(z^{k}),R(z):=z\cdot V(z)^{k}).$
We show that for general such $V$, we have $S_{1}(P)\neq S_{1}(R)$. Write $V(z)=a_{n}z^{n}+\cdots+a_{1}z+a_{0}$ with $(a_{n},\dots,a_{0})\in\mathbb{C}^{\times}\times\mathbb{C}^{n-1}\times\mathbb{C}^{\times}$. It is easy to see that $S_{1}(R)$ is the multi-set
$\{0,\rho_{R}(0)=a_{0}^{k},\rho_{R}(w_{1})=1+kw_{1}\frac{V^{\prime}(w_{1})}{V(w_{1})},\dots,\rho_{R}(w_{nk})=1+kw_{nk}\frac{V^{\prime}(w_{nk})}{V(w_{nk})}\},$
where $V(z)^{k}=a(z-w_{1})\cdots(z-w_{nk})+1$. One can check that
$\prod_{j=1}^{nk}\left(a_{0}-1-k\,w_{j}\frac{V^{\prime}(w_{j})}{V(w_{j})}\right)$
is a nonzero polynomial in $a_{0},\dots,a_{n}$; denote it by $f(a_{0},\dots,a_{n})$. When $(a_{n},\dots,a_{0})$ does not lie on the union of the hypersurfaces $a_{0}=a_{0}^{k}$ and $f(a_{0},\dots,a_{n})=0$, we have $\rho_{P}(0)=a_{0}\notin S_{1}(R)$, hence $S_{1}(P)\neq S_{1}(R)$. Therefore, for $(a_{n},\dots,a_{0})$ in a Zariski-dense open subset of $\mathbb{C}^{\times}\times\mathbb{C}^{n-1}\times\mathbb{C}^{\times}$, we have $S_{1}(P)\neq S_{1}(R)$.
@@
Let $P(z)=z^{2}(z^{3}+1)$ and $Q(z)=z^{2}(z+1)^{3}$, which satisfy the hypothesis of Theorem 4.6. Indeed, by Remark 4.7, for every odd positive integer 2n12n-1 ($n\in\mathbb{Z}_{>0}$), we have
@@
Using computer computations, one finds that $S_{2}(P)\neq S_{2}(Q)$ and $S_{4}(P)\neq S_{4}(Q)$, so $P$ and $Q$ do not have the same multiplier spectrum.
@@
Consider the Ritt move $(P(z)=z(z^{2}-3),R(z)=z(z-3)^{2})$. Then P(z2)+2=R(z)P(z-2)+2=R(z), hence $P$ and $R$ are conjugate and have the same multiplier spectrum.
Beyond the coincidence of the multiplier spectra on an infinite arithmetic progression under additional hypotheses, we can say more about the relation of the multiplier spectra in case (2) of Theorem 1.9.
@@
$P(z)=z^{r}\cdot R(z^{k})\quad\text{and}\quad Q(z)=z^{r}\cdot R(z)^{k}.$
@@
    For every $z_{0}\in{\rm Per}(P)$, there exist $w_{0}\in{\rm Per}(Q)$ and $m\in\mathbb{Z}_{>0}$ such that $\rho_{P}(z_{0})=\rho_{Q}(w_{0})^{m}$ and $n_{Q}(w_{0})\mid n_{P}(z_{0})$.
@@
    For every $w_{0}\in{\rm Per}(Q)$, there exist $z_{0}\in{\rm Per}(P)$ and $m\in\mathbb{Z}_{>0}$ such that $\rho_{Q}(w_{0})=\rho_{P}(z_{0})^{m}$ and $n_{Q}(w_{0})\mid n_{P}(z_{0})$.
In fact, in both (1) and (2) we can take $m=n_{P}(z_{0})/n_{Q}(w_{0})$.
@@
  1. (1)
    Let $z_{0}\in{\rm Per}(P)$ with exact period $l$. If $\rho_{P}(z_{0})=0$, take $w_{0}=\infty$ and $m=1$. Assume $\rho_{P}(z_{0})\neq 0$. In particular, $z_{0}\notin\{0,\infty\}$ since $r\geq 2$. From (4.11) (which holds without assuming $\gcd(r(r^{M}-1),k)=1$), we have n_Q(z_0^k)∣l and ρ_Q^∘l(z_0^k)=ρ_P^∘l(z_0)=ρ_P(z_0). Set $w_{0}=z_{0}^{k}$ and $m=l/n_{Q}(w_{0})$.
  2. (2)
    Let $w_{0}\in{\rm Per}(Q)$ with exact period $l$. If $\rho_{Q}(w_{0})=0$, take $z_{0}=\infty$ and $n=m=1$. Assume that $\rho_{Q}(w_{0})\neq 0$. In particular, $w_{0}\notin\{0,\infty\}$. By [FG22, Lemma 3.42], there exists $z_{0}\in{\rm PrePer}(P)$ such that $z_{0}^{k}=w_{0}$. Replacing $z_{0}$ by $P^{\circ l}(z_{0})$ for some suitable $l\geq 1$ if necessary, we may assume $z_{0}\in{\rm Per}(P)$. From (4.11), the choices $z_{0}$ and $m=n_{P}(z_{0})/l$ satisfy the requirement.
@@
We can give more explicit descriptions of $m$ and relations between $(S_{n}(P))$ and $(S_{n}(Q))$. For example, in the proof of 2, we must have
$P^{\circ jn}(z_{0})=\zeta_{j}z_{0}\quad(j=0,\dots,m-1),$
where $1=\zeta_{0},\dots,\zeta_{m-1}$ are distinct $k$-th roots of unity. In particular, $m\leq k$. We can deduce some other restrictions involving $\zeta_{j}$. Similarly, we have $m\leq k$ in 1.
Proposition 4.11 shows that in the Ritt move case, the multipliers of $P$ and $Q$ are mutually multiplicatively dependent. It is natural to consider multipliers modulo multiplicative dependence. For every rational map $f\in\mathbb{C}(z)$ of degree at least $2$, let $V(f)$ denote the $\mathbb{Q}$-vector space generated by $\{\chi_{f}(z):z\in{\rm Per}^{*}(f)\}\subseteq\mathbb{R}$, which is studied by the author, Ji, and Xie in [JXZ25].
@@
$P(z)=z^{r}\cdot R(z^{k})\quad\text{and}\quad Q(z)=z^{r}\cdot R(z)^{k}.$
@@
    $P$ is exceptional if and only if $Q$ is exceptional.
@@
(2) This follows from (1) and [JXZ25, Theorem 1.5]. (The converse of [JXZ25, Theorem 1.5] holds; see [Mil06].)
@@
Part (2) of Corollary 4.13 is a direct corollary of [Pak17, Theorem 4.4] (see also [FG22, Theorem 3.39]), which states: for two polynomials $f,g\in\mathbb{C}[z]$ of degree $d\geq 2$ with $\geq g, ff$ is conjugate to $z^{d}$ if and only if $g$ is conjugate to $z^{d}$, and $f$ is conjugate to $\pm T_{d}$ if and only if $g$ is conjugate to $\pm T_{d}$. Here $T_{d}(z)$ is the Chebyshev polynomial of degree $d$, and the exceptional polynomials of degree $d$ are exactly $z^{d},\pm T_{d}(z)$ up to conjugacy.
@@
Let $f,g\in\mathbb{C}(z)$ be two rational maps of the same degree $d\geq 2$. Assume first that $f\geq g$. It is easy to see that $V(f)\cap V(g)+W=V(f)+V(g)$ in $\mathbb{R}$ for some finite-dimensional subspace $W$ of $V(f)+V(g)$. Consequently, for intertwined $f$ and $g$, we have $V(f)\cap V(g)+W=V(f)+V(g)$ in $\mathbb{R}$ for some finite-dimensional subspace $W$ of $V(f)+V(g)$. Combined with [JXZ25, Theorem 1.3], we obtain that $f$ is exceptional if and only if $g$ is.

### v2

@@ 4.3. Ritt moves
(4.14) $Q^{\circ n}(z^{k})-z^{k}=P^{\circ n}(z)^{k}-z^{k}=(P^{\circ n}(z)-z)\prod_{j=1}^{k-1}(P^{\circ n}(z)-\zeta_{k}^{j}z).$
The left-hand side of (4.14) can be regarded as a polynomial in the variable $z^{k}-z_{0}^{k}$, and the right-hand side as a polynomial in $z-z_{0}$. Since $z_{0}\neq 0$, we have $\mathrm{ord}_{z-z_{0}}(z^{k}-z_{0}^{k})=1$ and $\mathrm{ord}_{z-z_{0}}(P^{\circ n}(z)-\zeta_{k}^{j}z)=0$ for $1\leq j\leq k-1$. Applying $\mathrm{ord}$ to both sides of (4.14), we obtain
$\mathrm{ord}_{z-z_{0}^{k}}(Q^{\circ n}(z)-z)$ $=\mathrm{ord}_{z^{k}-z_{0}^{k}}(Q^{\circ n}(z^{k})-z^{k})$
$=\mathrm{ord}_{z^{k}-z_{0}^{k}}(Q^{\circ n}(z^{k})-z^{k})\cdot\mathrm{ord}_{z-z_{0}}(z^{k}-z_{0}^{k})$
$=\mathrm{ord}_{z-z_{0}}(Q^{\circ n}(z^{k})-z^{k})$
$=\mathrm{ord}_{z-z_{0}}(P^{\circ n}(z)^{k}-z^{k})$
$=\mathrm{ord}_{z-z_{0}}(P^{\circ n}(z)-z).$
Thus $\mathrm{multi}_{F_{1}}(z_{0})=\mathrm{multi}_{F_{2}}(z_{0}^{k})\geq 2$ and (4.12) holds.
For the point $0=h(0)$, since $r\geq 2$, 0 is a superattracting fixed point of both $P^{\circ n}$ and $Q^{\circ n}$, i.e., $\rho_{P^{\circ n}}(0)=\rho_{Q^{\circ n}}(0)=0$. Hence $\mathrm{multi}_{F_{1}}(0)=\mathrm{multi}_{F_{2}}(0)=1$, concluding the proof of (4.12).
By (4.12) and $F_{1}=\#F_{2}$, if $h(z_{0})=h(z_{1})$ (in $\mathbb{C}$) implies $z_{0}=z_{1}$ (in $\mathbb{C}$) for $z_{0},z_{1}\in F_{1}$, then $S_{n}(P)=S_{n}(Q)$. In this case, we say that the integer $n$ is a good integer.
Claim: for every integer $n\geq 1$, either $n$ is good, or $\gcd(r^{n}-1,k)>1$.
Proof of the Claim: Assume $n\in\mathbb{Z}_{>0}$ is not good. Then there exist $z_{0}\neq z_{1}$ in $F_{1}$ with $z_{1}^{k}=z_{0}^{k}$. Hence $z_{1}\neq 0$ and $z_{1}=\zeta z_{0}$ for some $\zeta\in\mathbb{C}$ with $\zeta^{k}=1$ and $\zeta\neq 1$. It is easy to see that there exists $R_{n}(z)\in\mathbb{C}[z]$ with $R_{n}(0)\neq 0$ such that $P^{\circ n}(z)=z^{r^{n}}R_{n}(z^{k})$ and $Q^{\circ n}(z)=z^{r^{n}}R_{n}(z)^{k}$. From $P^{\circ n}(z_{0})=z_{0}\neq 0$ and $P^{\circ n}(\zeta z_{0})=\zeta z_{0}\neq 0$, we deduce $\zeta^{r^{n}}=\zeta$, so
$1<\mathrm{ord}_{\mathbb{C}^{\times}}(\zeta)\mid\gcd(r^{n}-1,k).$
Let $M$ be the minimal positive integer such that $\gcd(r(r^{M}-1),k)=1$. In particular, $r$ and $k$ are coprime. By Euler’s theorem, $r^{\varphi(k)}\equiv 1\ (\mathrm{mod}\ k)$. Let $d$ be the minimal positive integer with $r^{d}\equiv 1\ (\mathrm{mod}\ k)$. Then $d\mid\varphi(k)$. For every integer $N\geq 0$, we have
$r^{M+Nd}-1\equiv r^{M}-1\ (\mathrm{mod}\ k),$
so $\gcd(r^{M+Nd}-1,k)=1$. By the above Claim, every integer in the arithmetic progression $(M+Nd)_{N=0}^{\infty}$ is good. Since the sequence $(r^{n}\bmod k)_{n}$ has exact period $d$, $M\leq d$ by minimality. If $M=d$, then
$k\mid\gcd(r^{d}-1,k)=\gcd(r^{M}-1,k)=1,$
contradicting the assumption that $k\geq 2$. Thus $M<d$
@@
Let $k\geq 2$ be an integer, and $V(z)\in\mathbb{C}[z]$ a polynomial of degree $n\geq 1$ with $V(0)\neq 0$. Consider the Ritt move
$(P(z):=z\cdot V(z^{k}),Q(z):=z\cdot V(z)^{k}).$
We show that $S_{1}(P)\neq S_{1}(Q)$ for general such $V$. Write
$V(z)=a_{n}z^{n}+\cdots+a_{1}z+a_{0}\text{, where }(a_{n},\dots,a_{0})\in\mathbb{C}^{\times}\times\mathbb{C}^{n-1}\times\mathbb{C}^{\times}.$
It is easy to compute that $S_{1}(Q)$ is the multiset
$\left\{0,\rho_{Q}(0)=a_{0}^{k},\rho_{Q}(w_{1})=1+kw_{1}\frac{V^{\prime}(w_{1})}{V(w_{1})},\dots,\rho_{Q}(w_{nk})=1+kw_{nk}\frac{V^{\prime}(w_{nk})}{V(w_{nk})}\right\}$
of cardinality $nk+2$, where we write $V(z)^{k}=1+a_{n}^{k}(z-w_{1})\cdots(z-w_{nk})$. One can check that
$\prod_{j=1}^{nk}\left(a_{0}-1-k\,w_{j}\frac{V^{\prime}(w_{j})}{V(w_{j})}\right)=\prod_{j=1}^{nk}\left(a_{0}-1-k\,w_{j}V(w_{j})^{k-1}V^{\prime}(w_{j})\right)$
is a nonzero polynomial in $a_{0},\dots,a_{n}$; denote it by $f(a_{0},\dots,a_{n})$.
When $(a_{n},\dots,a_{0})$ does not lie on the union of the hypersurfaces $a_{0}=a_{0}^{k}$ and $f(a_{0},\dots,a_{n})=0$ in $\left(\mathbb{A}^{1}_{\mathbb{C}}\setminus\{0\}\right)\times\mathbb{A}_{\mathbb{C}}^{n-1}\times\left(\mathbb{A}^{1}_{\mathbb{C}}\setminus\{0\}\right)$, we have $\rho_{P}(0)=a_{0}\notin S_{1}(Q)$, hence $S_{1}(P)\neq S_{1}(Q)$. Therefore, for $(a_{n},\dots,a_{0})$ in a Zariski-dense open subset of $\left(\mathbb{A}^{1}_{\mathbb{C}}\setminus\{0\}\right)\times\mathbb{A}_{\mathbb{C}}^{n-1}\times\left(\mathbb{A}^{1}_{\mathbb{C}}\setminus\{0\}\right)$, we have $S_{1}(P)\neq S_{1}(Q)$.
@@
Let $P(z)=z^{2}(z^{3}+1)$ and $Q(z)=z^{2}(z+1)^{3}$, which satisfy the hypothesis of Theorem 4.6. Indeed, by Remark 4.7, for every $n\in\mathbb{Z}_{>0}$, we have
@@
Using computer computations, one finds that $S_{2}(P)\neq S_{2}(Q)$ and $S_{4}(P)\neq S_{4}(Q)$, so $P$ and $Q$ are not isospectral.
@@
Consider the Ritt move $(P(z)=z(z^{2}-3),Q(z)=z(z-3)^{2})$. Then P(z2)+2=Q(z)P(z-2)+2=Q(z), hence $P$ and $Q$ are conjugate and isospectral.
Beyond the coincidence of the multiplier spectra over an infinite arithmetic progression under additional hypotheses, we can consider the relation of the multiplier spectra of Ritt moves up to multiplicative dependence:
@@
$P(z)=z^{r}R(z^{k})\quad\text{and}\quad Q(z)=z^{r}R(z)^{k}.$
@@
    For every $z_{0}\in\mathrm{Per}(P)$, there exist $w_{0}\in\mathrm{Per}(Q)$ and $m\in\mathbb{Z}_{>0}$ such that $\rho_{P}(z_{0})=\rho_{Q}(w_{0})^{m}$ and $n_{Q}(w_{0})\mid n_{P}(z_{0})$.
@@
    For every $w_{0}\in\mathrm{Per}(Q)$, there exist $z_{0}\in\mathrm{Per}(P)$ and $m\in\mathbb{Z}_{>0}$ such that $\rho_{Q}(w_{0})=\rho_{P}(z_{0})^{m}$ and $n_{Q}(w_{0})\mid n_{P}(z_{0})$.
In fact, for both (1) and (2) we can choose $m=n_{P}(z_{0})/n_{Q}(w_{0})$.
@@
(1) Let $z_{0}\in\mathrm{Per}(P)$ with exact period $n$. If $\rho_{P}(z_{0})=0$, then take $w_{0}=\infty$ and any $m\in\mathbb{Z}_{>0}$. Assume $\rho_{P}(z_{0})\neq 0$. In particular, $z_{0}\notin\{0,\infty\}$ since $r\geq 2$. From (4.13) (which holds without assuming $\gcd(r(r^{M}-1),k)=1$), we have
$n_{Q}(z_{0}^{k})\mid n\quad\text{and}\quad\rho_{Q^{\circ n}}(z_{0}^{k})=\rho_{P^{\circ n}}(z_{0})=\rho_{P}(z_{0}).$
Set $w_{0}=z_{0}^{k}$ and $m=n/n_{Q}(w_{0})$.
(2) Let $w_{0}\in\mathrm{Per}(Q)$ with exact period $n$. If $\rho_{Q}(w_{0})=0$, take $z_{0}=\infty$ and any $m\in\mathbb{Z}_{>0}$. Assume $\rho_{Q}(w_{0})\neq 0$. In particular, $w_{0}\notin\{0,\infty\}$. By [FG22, Lemma 3.42], there exists $z_{0}\in\mathrm{PrePer}(P)$ such that $z_{0}^{k}=w_{0}$. Replacing $z_{0}$ by $P^{\circ l}(z_{0})$ for some suitable $l\in\mathbb{Z}_{>0}$ if necessary, we may assume that $z_{0}\in\mathrm{Per}(P)$. It is clear that $n=n_{Q}(w_{0})\mid n_{P}(z_{0})$. By (4.13), the choices $z_{0}$ and $m=n_{P}(z_{0})/n$ satisfy the requirement. ∎
@@
We can give more explicit descriptions of $m$ and relations between the multiplier spectra $(S_{n}(P))$ and $(S_{n}(Q))$. For example, in the proof of the $\rho_{Q}(w_{0})\neq 0$ case of (2), we must have
$P^{\circ jn}(z_{0})=\zeta_{j}z_{0}\quad(0\leq j\leq m-1),$
where $1=\zeta_{0},\dots,\zeta_{m-1}$ are distinct $k$-th roots of unity. In particular,
$m=\frac{n_{P}(z_{0})}{n_{Q}(w_{0})}\leq k.$
We could also deduce some other restrictions involving $\zeta_{j}$. Similarly, we can require that $m=\frac{n_{P}(z_{0})}{n_{Q}(w_{0})}\leq k$ in (1).
Proposition 4.11 shows that for the Ritt move $(z^{r}R(z^{k}),z^{r}R(z)^{k})$ with $r\geq 2$, the multipliers of $P$ and $Q$ are mutually multiplicatively dependent. The author has studied multipliers modulo multiplicative dependence in the joint work [JXZ25] with Ji and Xie. For every rational map $f\in\mathbb{C}(z)$ of degree at least $2$, let $V(f)\subseteq\mathbb{R}$ be the $\mathbb{Q}$-vector space generated by $\{\chi_{f}(z)\mid z\in\mathrm{Per}^{*}(f)\}$.
@@
$P(z)=z^{r}R(z^{k})\quad\text{and}\quad Q(z)=z^{r}R(z)^{k}.$
@@
    $P$ is exceptional if and only if $Q$ is exceptional;
@@
(2) This follows from (1) and [JXZ25, Theorem 1.5]. (Note that the converse of [JXZ25, Theorem 1.5] also holds; see [Mil06b].)
@@
Part (2) of Corollary 4.13 is a direct corollary of [Pak17, Theorem 4.4] (see also [FG22, Theorem 3.39]), which asserts that for two polynomials $f,g\in\mathbb{C}[z]$ of degree $d\geq 2$ with $f\geq g$, the following statements hold:
  • $f$ is conjugate to $z^{d}$ if and only if $g$ is conjugate to $z^{d}$;
  • $f$ is conjugate to $\pm T_{d}$ if and only if $g$ is conjugate to $\pm T_{d}$.
Recall that up to conjugacy, exceptional polynomials of degree $d$ are exactly $z^{d}$ and $\pm T_{d}(z)$, and that $T_{d}$ is conjugate to $T_{d}$ if and only if $d$ is even.
@@
Let $f,g\in\mathbb{C}(z)$ be two rational maps of the same degree $d\geq 2$. Assume first that $f\geq g$. Arguing similarly as in the proof of Proposition 4.11, it is easy to see that
(4.15) $V(f)\cap V(g)+W=V(f)+V(g)$
for some finite-dimensional subspace $W$ of $V(f)+V(g)$. Consequently, for intertwined $f$ and $g$, this conclusion (4.15) still holds. Combined with [JXZ25, Theorem 1.4] (whose converse also holds [Mil06b]), we obtain a proof of the fact that for intertwined $f$ and $g$, $f$ is exceptional if and only if $g$ is exceptional.